OpenStudy (anonymous):

log[log(x+4)]=0 what is value of x and how

OpenStudy (kirbykirby):

Say \(\log(x+4)=y\), then \(\log y = 0 \implies e^0 = y \implies y = 1\) Thus, \(\log(x+4)=1\), so \(e^1=x+4=\implies x=e-4\)

OpenStudy (kirbykirby):

Oh wait actually, are you using log in base \(e\) or 10?

OpenStudy (anonymous):

thanks bro i understand

OpenStudy (anonymous):

log[log(x+4)-4]=0 then x=-4 how

OpenStudy (kirbykirby):

o_O If you're in base \(e\), you should get \(e^5-4\). If you are in base 10, you should get \(10^5-4\).

OpenStudy (kirbykirby):

In fact, if x=-4 here, you get \(\log(\log(-4+4)-4)=\log(\log(0)-4)\) and \(\log0\) is undefined